# FDSG: Forecasting Dynamic Scene Graphs

### v1

@@ C.3 More Ablation Study
C.4 Entity Detection
We use Object Detection Recall to assess how well predicted entities match the ground truth. We evaluate under two settings: (1) SGA where IoU $\geq 0$, i.e. only evaluates label accuracy; and (2) SGF where IoU $\geq 0.5$, i.e. includes bounding box quality. We compare against a Baseline using the predicted scene graph from the last observation and an Oracle using the last ground truth. We also report Average Precision (AP) with IoU $\geq 0.5$ in Table 15. Our model clearly outperforms the Baseline, in particular in the SGA setting, showing promising capability to discover novel nodes. Our model sometimes closely trails the Oracle due to minor prediction errors for the last observation, and in the SGF setting due to accumulated error in label and bounding box prediction.
Table 15: Object detection recall and average precision for SGF with different observation fraction $\mathcal{F}$.
SGA (IoU $\geq$ 0) SGF (IoU $\geq$ 0.5)
Baseline Oracle Ours Baseline Oracle Ours Baseline Oracle Ours
0.3 77.00 77.05 81.57 37.07 39.23 40.67 9.1 14.0 10.8
0.5 80.11 82.75 83.36 40.33 44.01 43.85 10.5 16.3 12.1
0.7 84.44 89.76 87.38 45.67 51.27 47.87 13.3 23.7 16.7
0.9 86.99 96.01 91.89 55.25 64.95 58.51 18.7 37.5 22.2
Appendix D Qualitative Results
D.1 Attention Maps
@@
Figure 5: Visualization of sampling points of the Deformable Attention in the last Cross-Attention layer of FDSG’s entity decoder for the DSGG task. Brighter color indicates a larger attention weight associated with the sampling point. We also label the detected triplets and the bounding boxes.
In Fig. 5, we visualize the sampling points of the Deformable Attention [50] in our FDSG’s entity decoder. The visualized sampling points are from the last Cross-Attention layer of the entity decoder. We show the top-5 most confident queries in terms of the overall classification score of entities and predicate. There are 128 sampling points associated with each query, and one attention weight associated with each sampling point. The larger attention weights are visualized in brighter color (e.g. yellow) while the smaller weights are shown in darker color (e.g. dark blue). Interestingly, the deformable attention not only attends to the entities, but also attempts to locate the interaction between the subject and the object. This is especially obvious in the first two columns, where the corresponding predicate is looking / not looking at, where the Attention mechanism concentrates around the direction where the central person is looking. We can also see that this direction is actually blocked by the table, which is possibly the reason that the model mistakenly predicts person – not looking at – bottle.
D.2 More DSGG and SGF Results
@@
In Figs. 6-9, we present more DSGG and SGF results from FDSG. In particular, the results shown in Fig. 6 are contrasted with the qualitative results in Sec. 4.4 in our main paper. In this case, for SGF, the observed video fraction is $\mathcal{F}=0.5$. In comparison, the Sec. 4.4 example has $\mathcal{F}=0.7$, and in the last observed frame, the bottle on the left is clearly visible, and the person is also looking to the left, showing a clear intention to interact with the bottle. However, when $\mathcal{F}=0.5$, FDSG fails to detect the bottle due to partial occlusion in the last observed frame. In addition, the person is looking at the notebook and has not yet started the interaction with the bottle. Consequently, in SGF, our model does not predict any interaction between the person and the bottle. Instead, FDSG forecasts food / sandwich showing up in the center, again demonstrating that it has a comprehensive understanding of the scene and is capable of forecasting unseen actions.
In Figs. 7-9, for SGF, the observed video fraction is $\mathcal{F}=0.5$. In Fig. 7, the video shows a man cleaning his room with a broom. FDSG is able to forecast the movement of the broom. In Fig. 8, the video shows a man leaving his laptop to fetch some food. FDSG forecasts that the man would interact with the box in front of him, instead of the food that is farther away. In Fig. 9, the video shows a woman taking off her shoes. FDSG forecasts that the woman would interact with her clothes, which is very close to the precise action of removing shoes.
Figure 6: Visualization of Dynamic Scene Graph Generation (DSGG) and Scene Graph Forecasting (SGF) outputs from our FDSG. The first row shows frames and ground truth / predicted entity bounding boxes, while the second row shows corresponding scene graphs. In the scene graphs, incorrect predicates are highlighted with text in red, while entities / predicates with missing labels in the ground truth are highlighted in orange. For (a) observed frames, we provide our DSGG results and compare them with the ground truth. For (b) unobserved frames, we contrast our SGF results to the ground truth.
@@
Figure 7: Visualization of Dynamic Scene Graph Generation (DSGG) and Scene Graph Forecasting (SGF) outputs from our FDSG. The first two rows show frames with ground truth entity bounding boxes and ground truth scene graphs, while the last two rows show generated scene graphs from DSGG or SGF. In the scene graphs, incorrect predicates are highlighted with text in red, while entities / predicates with missing labels in the ground truth are highlighted in orange.
@@
Figure 8: Visualization of Dynamic Scene Graph Generation (DSGG) and Scene Graph Forecasting (SGF) outputs from our FDSG. The first two rows show frames with ground truth entity bounding boxes and ground truth scene graphs, while the last two rows show generated scene graphs from DSGG or SGF. In the scene graphs, incorrect predicates are highlighted with text in red, while entities / predicates with missing labels in the ground truth are highlighted in orange.
@@
Figure 9: Visualization of Dynamic Scene Graph Generation (DSGG) and Scene Graph Forecasting (SGF) outputs from our FDSG. The first two rows show frames with ground truth entity bounding boxes and ground truth scene graphs, while the last two rows show generated scene graphs from DSGG or SGF. In the scene graphs, incorrect predicates are highlighted with text in red, while entities / predicates with missing labels in the ground truth are highlighted in orange.

### v2

@@ 5.3.2 Realization of the LDM
Figure 4: Visualization of bounding boxes and generated scene graphs from DSGG (left) and SGF (right), under observation fractions: top row - $\mathcal{F}=0.5$, bottom row - $\mathcal{F}=0.7$.
We evaluate different instantiations of the Location Dynamics Model (LDM) from Eq. (7) with results summarized in Table XI. When using residual updates for entity location representations (Eq. (9)), the Triplet Dynamics Model (TDM) also predicts these residuals. We hypothesize that this coupling helps TDM better model predicate dynamics when they are linked to entity motion, which is supported by the observed performance trends. Among the variants, Identity Mapping offers the best overall performance: top in SGA, second-best in DSGG and SGF, with only marginal gaps from the top models in the latter two tasks. The 3-layer MLP achieves the best DSGG performance but underperforms in SGA. NeuralSDE excels in SGF, indicating stronger entity forecasting, but lags in DSGG and SGA, suggesting reduced effectiveness in predicate anticipation. Based on overall performance, simplicity, and compactness, Identity Mapping is adopted.
5.3.3 Comparison of Aggregation Strategy
In FDSG’s Temporal Aggregation Module, we first aggregate information from reference frames with lower uncertainty, followed by forecasts with higher uncertainty. In Table XII, we test two variants: (a) aggregating forecasts before observations, and (b) aggregating both information in a single cross-attention layer. Neither outperforms our current design in FDSG.
5.4 Qualitative Results
5.4.1 Forecast with Different Observation Fraction
@@
Figure 5: Visualization of sampling points of the Deformable Attention in the last Cross-Attention layer of FDSG’s entity decoder for the DSGG task. A brighter color indicates a larger attention weight associated with the sampling point. We also label the detected triplets and the bounding boxes.
@@
In Fig. 4, we present DSGG and SGF results from FDSG under different observation fractions $\mathcal{F}$. In particular, the results shown in the top row use an observed video fraction $\mathcal{F}=0.5$, while the bottom row uses $\mathcal{F}=0.7$. When $\mathcal{F}=0.5$, FDSG fails to detect the bottle on the left due to partial occlusion in the last observed frame. In addition, the person is looking at the notebook and has not yet started the interaction with the bottle. Consequently, in SGF, our model does not predict any interaction between the person and the bottle. Instead, FDSG forecasts food/sandwich showing up in the center, demonstrating a comprehensive understanding of the scene and its capability to forecast unseen entities. In comparison, the bottom row uses $\mathcal{F}=0.7$, where in the last observed frame the bottle is clearly visible, and the person is also looking to the left, showing a clear intention to interact with the bottle. As a result, in SGF, our FDSG successfully predicts the movement of the bottle from the left to the center. In addition to the moving object (the bottle), FDSG also correctly predicts the location of static objects (the person, table, and notebook), demonstrating the effectiveness of our proposed Location Dynamics Model using Identity Mapping for the query locations.
5.4.2 Attention Maps
@@
Figure 6: Visualization of Dynamic Scene Graph Generation (DSGG) and Scene Graph Forecasting (SGF) outputs from our FDSG. The first two rows show frames with ground truth entity bounding boxes and ground truth scene graphs, while the last two rows show generated scene graphs from DSGG (left) and SGF (right). In the scene graphs, incorrect predicates are highlighted with text in red, while entities/predicates with missing labels in the ground truth are highlighted in orange.
In Fig. 5, we visualize the sampling points of the Deformable Attention module [50] within the entity decoder of our FDSG model. The visualizations are taken from the final cross-attention layer and correspond to the top five most confident queries, ranked by the overall classification scores of entities and predicates. Each query is associated with 128 sampling points, each assigned an attention weight. Brighter colors (e.g., yellow) indicate higher attention weights, while darker colors (e.g., dark blue) represent lower ones.
Notably, the deformable attention mechanism attends not only to the entities themselves but also to the regions representing interactions between subjects and objects. This is particularly evident in the first two columns, where the predicted predicates are looking at and not looking at. In these cases, attention focuses around the direction in which the central person is facing. Interestingly, this direction is obstructed by a table, which may explain the model’s misprediction: person – not looking at – bottle.
5.4.3 More DSGG and SGF Results
@@
Figure 7: Visualization of Dynamic Scene Graph Generation (DSGG) and Scene Graph Forecasting (SGF) outputs from our FDSG. The first two rows show frames with ground truth entity bounding boxes and ground truth scene graphs, while the last two rows show generated scene graphs from DSGG (left) and SGF (right). In the scene graphs, incorrect predicates are highlighted with text in red, while entities/predicates with missing labels in the ground truth are highlighted in orange.
@@
Figure 8: Visualization of Dynamic Scene Graph Generation (DSGG) and Scene Graph Forecasting (SGF) outputs from our FDSG. The first two rows show frames with ground truth entity bounding boxes and ground truth scene graphs, while the last two rows show generated scene graphs from DSGG (left) and SGF (right). In the scene graphs, incorrect predicates are highlighted with text in red, while entities/predicates with missing labels in the ground truth are highlighted in orange.
Figs. 68 showcase additional DSGG and SGF results produced by FDSG. For the SGF task, the observed video fraction is set to $\mathcal{F}=0.5$. In Fig. 6, the video depicts a man cleaning his room with a broom. In the DSGG results, FDSG successfully detects the presence of clothes, even though they are not included in the ground truth annotations. For the SGF task, the model predicts the dynamics of the broom’s bounding box. Fig. 7 shows a man leaving his laptop to get food. FDSG predicts an increase in the height of the man’s bounding box, correctly inferring that he is standing up. It also forecasts an interaction with a nearby box rather than the food located farther away, which is not labeled in the observed frames. In Fig. 8, the video shows a woman taking off her shoes. FDSG predicts an interaction with her clothes as shoes are not visible in the observed part, which is still closely aligned with the action of removing shoes.
6 Conclusion